# Consumption Partial Insurance

### v5

@@ B.3 Quadratic consumption function parameters
Another equation that pins down $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ is $2\gamma_{\zeta_{t}}\phi^{(1)}_{t}+2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}))\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$, which, abstracting from income measurement error, corresponds to the second line in (B.1). Absent transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and squared shocks so it captures the pass-through of shocks of large magnitude. We highlight again three cases (for $\phi^{(1)}_{t}$ fixed):
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}=0$. This identifies $\phi^{(2)}_{t}=0$, so tail shocks transmit similarly to non-tail ones. In fact, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})=2\gamma_{\zeta_{t}}\phi^{(1)}_{t}$ is directly implied by $\Delta^{2}c_{it}=\phi_{t}^{(1)}(\zeta_{it}+\zeta_{it-1})$, so this configuration identifies the linear consumption process.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}<0$, which, for $\gamma_{\zeta_{t}}<0$, is true when ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$ is negative and large. This suggests that large shocks (of which negative shocks are the most likely) transmit strongly with a negative coefficient. It thus identifies $\phi^{(2)}_{t}<0$. Ceteris paribus, a less leptokurtic income distribution (lower $\kappa_{\zeta_{t}}$) turns $\phi^{(2)}_{t}$ more negative; a bell-shaped distribution is associated with fewer large shocks, thus necessitating a stronger transmission of said shocks to maintain a given excess co-movement.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}>0$. This identifies $\phi^{(2)}_{t}>0$ for analogous reasons. For given excess co-movement, the smaller the kurtosis of permanent shocks, the larger the value of $\phi^{(2)}_{t}$ must be to achieve said excess co-movement.
@@
An interesting special case arises when the distributions of permanent and transitory shocks are normal, have the same variance between them, and there is no measurement error. In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the hypothetical distribution. Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$. Identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks. Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks. Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
@@ B.4 Targeted moments
There is a large number of over-identifying moments that we target in the estimation of the various specifications of the model. Table B.1 lists the income moments targeted in each case, while table B.2 lists the moments targeted in the different specifications of the consumption model. For simplicity, we illustrate the next statements assuming stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters. These are not identifying assumptions but they make the illustration more compact by allowing us to pool common terms together.
@@
1 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$
2 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$
3 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$
4 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$
5 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$
6 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})$
7 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it-2})^{2})$
8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}y_{it-2})$
9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{3})$
Notes: The table lists the moments targeted in the estimation of the various specifications of the income process. $\Delta^{2}y_{t}=\Delta y_{t}+\Delta y_{t-1}=y_{t}-y_{t-2}$ indicates a first difference of $y$ between periods $t$ and $t-2$.
@@ B.4.1 Moments of income
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ $=2\sigma_{\zeta}^{2}+2\sigma_{v}^{2}+2\sigma_{{u}^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ $=-\sigma_{v}^{2}-\sigma_{{u}^{y}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$ $=2\gamma_{\zeta}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$ $=\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$ $=-\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})$ $=2\kappa_{\zeta}+2\kappa_{v}+6(\sigma_{\zeta}^{2})^{2}+6(\sigma_{v}^{2})^{2}+12(\sigma_{u^{y}}^{2})^{2}+24\sigma_{\zeta}^{2}\sigma_{v}^{2}+24\sigma_{v}^{2}\sigma_{u^{y}}^{2}+24\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it-2})^{2})$ $=\kappa_{v}+4(\sigma_{\zeta}^{2})^{2}+3(\sigma_{v}^{2})^{2}+6(\sigma_{u^{y}}^{2})^{2}+8\sigma_{\zeta}^{2}\sigma_{v}^{2}+8\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}y_{it-2})$ $=-\kappa_{v}-3(\sigma_{v}^{2})^{2}-6(\sigma_{u^{y}}^{2})^{2}-6\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{3})$ $=-\kappa_{v}-3(\sigma_{v}^{2})^{2}-6(\sigma_{u^{y}}^{2})^{2}-6\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
@@
3 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$
4 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$
@@
8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$
9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$
10 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$
11 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta c_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{2})$
12 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$
13 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$
14 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$
15 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$

### v6

@@ B.3 Quadratic consumption function parameters
Another equation that pins down $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ is $2\gamma_{\zeta_{t}}\phi^{(1)}_{t}+2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}))\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})$, which, abstracting from income measurement error, corresponds to the second line in (B.1). Absent transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and squared shocks so it captures the pass-through of shocks of large magnitude. We highlight again three cases (for $\phi^{(1)}_{t}$ fixed):
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}=0$. This identifies $\phi^{(2)}_{t}=0$, so tail shocks transmit similarly to non-tail ones. In fact, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})=2\gamma_{\zeta_{t}}\phi^{(1)}_{t}$ is implied by $\Delta^{2}c_{it}=\phi_{t}^{(1)}(\zeta_{it}+\rho\zeta_{it-1})$, so this configuration identifies the linear consumption process.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}<0$, which, for $\gamma_{\zeta_{t}}<0$, is true when ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})$ is negative and large. This suggests that large shocks (of which negative shocks are the most likely) transmit strongly with a negative coefficient. It thus identifies $\phi^{(2)}_{t}<0$. Ceteris paribus, a less leptokurtic income distribution (lower $\kappa_{\zeta_{t}}$) turns $\phi^{(2)}_{t}$ more negative; a bell-shaped distribution is associated with fewer large shocks, thus necessitating a stronger transmission of said shocks to maintain a given excess co-movement.
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}>0$. This identifies $\phi^{(2)}_{t}>0$ for analogous reasons. For given excess co-movement, the smaller the kurtosis of permanent shocks, the larger the value of $\phi^{(2)}_{t}$ must be to achieve said excess co-movement.
@@
An interesting special case arises when $\rho=1$, the distributions of permanent and transitory shocks are normal and have the same variance, and there is no measurement error. In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the hypothetical distribution. Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$. Identification does not require shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks. Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks. Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
@@ B.4 Targeted moments
The estimation targets several over-identifying moments. Table B.1 lists the target income moments, while table B.2 lists the moments targeted in the various specifications of the consumption model. For simplicity, we illustrate these moments assuming stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters. These are not identifying assumptions but they make the illustration more compact by allowing us to pool common terms together.
@@
1 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$
2 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$
3 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$
4 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3})$
5 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it-2})$
6 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$
7 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{4})$
8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$
9 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}_{\rho}y_{it-2})$
10 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{3})$
Notes: The table lists the moments targeted in the estimation of the various specifications of the income process. $\Delta^{2}y_{t}=y_{t}-y_{t-2}$ indicates a first difference between periods $t$ and t2t-2; $\Delta^{2}_{\rho}y_{t}=y_{t}-\rho^{2}y_{t-2}$ indicates the analogous quasi-difference. Note that $\Delta^{2}y_{t}=\Delta^{2}_{\rho}y_{t}$ if $\rho=1$ (baseline permanent-transitory process).
@@ B.4.1 Moments of income
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$ $=(1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})\sigma_{v}^{2}+(1+\rho^{4})\sigma_{{u}^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$ $=-\rho^{2}\sigma_{v}^{2}-\rho^{2}\sigma_{{u}^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$ $=0$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3})$ $=(1+\rho^{3})\gamma_{\zeta}+(1-\rho^{6})\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it-2})$ $=\rho^{4}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$ $=-\rho^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{4})$ $=(1+\rho^{4})\kappa_{\zeta}+(1+\rho^{8})\kappa_{v}+6\rho^{2}(\sigma_{\zeta}^{2})^{2}+6\rho^{4}(\sigma_{v}^{2})^{2}+3(1+\rho^{4})^{2}(\sigma_{u^{y}}^{2})^{2}$
$+6(1+\rho^{4})\Big((1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}\Big)$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$ $=\rho^{4}\kappa_{v}+(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+(1+\rho^{4}+\rho^{8})(\sigma_{v}^{2})^{2}+(1+4\rho^{4}+\rho^{8})(\sigma_{u^{y}}^{2})^{2}$
$+2(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{v}^{2}+2(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+(2(1+\rho^{4})^{2}+4\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}_{\rho}y_{it-2})$ $=-\rho^{6}\kappa_{v}-3\rho^{2}(\sigma_{v}^{2})^{2}-3\rho^{2}(1+\rho^{4})(\sigma_{u^{y}}^{2})^{2}$
$-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-6\rho^{2}(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{3})$ $=-\rho^{2}\kappa_{v}-3\rho^{6}(\sigma_{v}^{2})^{2}-3\rho^{2}(1+\rho^{4})(\sigma_{u^{y}}^{2})^{2}$
$-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-6\rho^{2}(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
@@
3 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$
4 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$
@@
8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$
9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{2})$
10 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$
11 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta c_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{2})$
12 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$
13 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$
14 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$
15 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$